OpenStudy (jenniferjuice):

Given that f(x) = 5x − 10 and g(x) = x + 3, solve for f(g(x)) when x = −1. −30 0 10 12

OpenStudy (kirbykirby):

plug in all of g(x) into f(x) wherever you see "x" f(g(x)) = 5(g(x)) - 10 = 5(x+3) - 10 Then just substitute that expression above with x=-1

OpenStudy (jenniferjuice):

so its 10 ?

OpenStudy (kirbykirby):

no

OpenStudy (kirbykirby):

make sure you calculate all of 5(x+3) - 10

OpenStudy (jenniferjuice):

oh no its 0 ?

OpenStudy (kirbykirby):

yes =]

OpenStudy (jenniferjuice):

thank you :D

OpenStudy (kirbykirby):

yw!